. If a proton and anti-proton come close to each other and annihilate, how much energy will be released ? 

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
The energy released = total energy of proton and anti-proton. 
Rest energy of a proton/anti-proton = 938.257 MeV (Mega-electronvolts) (using E = mc^2) 
2 * 938.257 = 1876.514 MeV (because there are 2 particles) 
1876.514 * 10^6 = 1876514000 eV (get from MeV to eV) 
1 eV = 1.60 * 10^-19J (convert to joules using charge of an electron (because we're in electron volts)) 
So 1876514000 eV = (1876514000 * (1.60 * 10^-19)) J 
= 3.00 * 10^-10 J 
Regards
Arun (askIITans forum expert)